April 12, 2017, Production & Analysis Working Group

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

The meeting will be on Wednesday April 12, 2017 at 2:00 pm EDT. For those people at Jefferson Lab, the meeting will be in room F326-327.

Meeting Connections

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/115815824.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 115815824.
  3. To join via phone, use one of the following numbers and the Conference ID: 115815824.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.

Reconstruction & Analysis Studies Status


  1. Announcements
  2. Collaboration Meeting May 15-17, 2017
  3. Data Production
  4. Analysis Focus
  5. This Week's Studies
  6. Any other studies
  7. Upcoming Study Schedule

This Week's Topics

  • Talks: Summary/overviews only. Detailed discussions should be held in the appropriate working groups.
  • Talks should be limited to 10 + 5 minutes.

Data Production

  1. Monitoring Update --- Thomas Britton
  2. Calibration Update --- Sean Dobbs
  3. Preparing for the first Spring 2017 production launch
  4. Processing Update --- Alex Austregesilo

Analysis Focus: Total Cross Sections

  1. J/Ψ, φ --- Lubomir
  2. η, ρ, ω, φ --- Simon
  3. φ --- Thomas
  4. Any others?

Studies: Alignment & Track/Shower Efficiencies

(i.e. Is the reconstruction working?) Updates on experiment, simulation, and the comparison between the two.

  1. Drift chamber alignment --- Mike Staib
  2. Tracking Efficiencies --- Paul Mattione
  3. BCAL Shower Efficiencies --- Elton Smith
  4. FCAL Shower Efficiencies --- Jon Zarling

Data Production Milestones

  1. Fri. April 21 - Finalize calibrations/software for low-rate running. Run final tests.
    • Produce Spring 2016 "gold runs" as soon as alignment is ready
  2. Fri. April 28 - Start production for low-rate running.
  3. Fri. May 5 - Finalize calibrations/software for high-rate running. Run final tests.
  4. Mon. May 8 -- Fri. May 12 (?) - Scicomp tape library down
  5. Fri. May 19 - Start production for high-rate running.

Upcoming Study Schedule

  • Updates on experiment, simulation, and the comparison between the two.
  • Summary/overviews only. Detailed discussions should be held in the appropriate working groups.
  • All talks should be limited to 10 + 5 minutes.
  • Next Week: Other
  1. Track / Hit Matching: BCAL, FCAL, TOF, SC --- Paul Mattione
  2. Efficiencies: TOF, SC
  3. Means & Resolutions (time, energy, dE/dx): Tracking, BCAL, FCAL, SC, TOF
  4. Uncertainties: PID (BCAL, FCAL, TOF, dE/dx), Kinfit (BCAL, FCAL, tracking)
  5. Channel/Analysis Studies: Branching ratios, cross sections, SDMEs, beam asymmetries
  6. Other reconstruction/analysis issues
  • Week after next: Beamline & Triggering
  1. Flux --- Justin
  2. Beam energy --- Beamline Group
  3. Polarization (TPOL & lineshape) --- Beamline Group
  4. Beam Asymmetries --- Alex Austregesilo
  5. Trigger emulation --- Alex Somov
  6. Triggering efficiency --- Alex Somov
  • The following week: Hit Efficiencies

(i.e. Is the detector working?) Updates on experiment, simulation, and the comparison between the two.

  1. CDC Hit Efficiencies --- Naomi Jarvis
  2. FDC Hit Efficiencies --- Alex Austregesilo
  3. BCAL Hit Efficiencies --- Elton Smith
  4. FCAL Hit Efficiencies --- Jon Zarling


Data Production

  1. Sean gave an update on the status of calibrations. The BCAL timing is being finalized, after that higher-level calibrations need to be checked. Aiming to be done by end of next week, but this work will likely drive the schedule. FCAL 4 ns timing shifts are few-percent on first glance, might just have to loosen selections for this first production run. Nathan is finalizing TAGH/PS calibrations, should be done early next week. Sean is checking TAGM/RF alignment with higher-statistics data needed for low-yield fibers. Drift chamber alignment is reported on below.
    • Expect full spring data set calibrations to be finished by end of month.
  2. Waiting on results for the current monitoring launch that is using the first 10 and last 10 files of each run - results delayed due to farm competition, should be ready tomorrow?

Analysis Focus: Total Cross Sections

  1. Simon is working on the robustness of his omega peak fitter, since it seems to be giving some strange results in certain energy bins.
  2. Thomas is working on his fitter as well, and expects to have some preliminary cross sections using efficiencies from geant3 and geant4 simulations separately.
  3. Cris is working on a final estimate of efficiencies for the two omega decays that he's looking at in the context of the B-boson search.

Studies: Alignment & Track/Shower Efficiencies

  1. Mike Staib reports that he has some preliminary alignment parameters that are ready to go. There are still some improvements that can be made, for example the FDC wire plane alignment implies a shrinking in the size of the FDC chambers, but this will be an improvement over what we currently have. He is doing some final code tests and will submit a pull request by the end of the week.
  2. Paul reports that he was looking more into his preliminary tracking efficiency calculations, and noticed that there is a long tail of tracks with large (>40%) difference in momentum from what was expected from the missing momentum in the rest of the event. He identified several events with delta p / p ~ 50 - 60%, and gave them to Simon to look at.
    • Matt suggested looking at simulations to improve the error estimates and event selections.
  3. Jon and Elton report no new news since last week